Give the precise thermochemical equation to which the following statements refer.
(i) The standard enthalpy of formation Of methane is - 74.8 K J/mol
(ii) The mean bond energy of C - H bond in methane is 413 KJ/mol
(iii) The enthalpy of atomization of carbon is 715 KJ/mol
1
Expert's answer
2020-12-06T10:25:55-0500

(i) C(solid, graphite) + 2H2(gas) = CH4(gas) "\Delta"fHo = -74 kJ/mol

(ii) 1/4CH4(gas) = 1/4C(gas) + H(gas) "\Delta"rH = 413 kJ/mol

(iii) C(solid, graphite) = C(gas) "\Delta"atomH = 715 kJ/mol